# Continuous Family of Conformal Field Theories and Exactly Marginal Operators

### v1

@@ I Introduction and Summary
If exactly marginal operators exist in a conformal field theory (CFT), they generate a conformal manifold, which parameterizes a continuous family of CFTs [Zamolodchikov1986]. One might naturally ask if the converse also holds: Does a conformal manifold require the existence of exactly marginal operators? In this letter, we address this question by deriving the existence of exactly marginal operators under certain assumptions.
@@
The existence of exactly marginal operators has important implications for the Swampland program, particularly for the Distance Conjecture [Ooguri:2006in]. The earliest distance conjecture may have been stated by Albert Einstein in his autobiographical notes, in which he claimed that there are no free parameters in the basic equations of physics. This was reformulated as Conjecture 0 in [Ooguri:2006in] as “the moduli space of a consistent quantum gravity is parameterized by inequivalent expectation values of massless scalar fields.” The existence of exactly marginal operators for a conformal manifold allows us to prove Conjecture 0 for gravitational theories in anti-de Sitter (AdS) space as follows [Ooguri:2022]: If there is a continuous family of gravitational theories in AdS, then there must be a corresponding family of CFTs. If we can show that each CFT parameter is associated with an exactly marginal operator, the AdS/CFT dictionary would tell us that the exactly marginal operator corresponds to a massless scalar field with a flat potential, and the asymptotic behavior of this field determines the parameters. Therefore, every parameter in quantum gravity in AdS is related to a value of a dynamical field, as claimed in Conjecture 0.111There has been an interesting development in interpreting the conjecture in AdS [Banks:2025nfe, Sen:2025bmj, Sen:2025ljz, Sen:2025oeq].
@@
We note an interesting analogy with the absence of global symmetry in quantum gravity in AdS, which was proven in [Harlow:2018jwu, Harlow:2018tng] using the splittability of symmetry generators [Buchholz:1985ii] in CFT. For continuous symmetry, the splittability is the consequence of the well-known fact that the symmetry generators are expressed as exponentiated integrals of the Noether currents over a Cauchy surface. If we regard a continuous deformation of CFT as (1)(-1)-form global symmetry, the existence of the corresponding exactly marginal operator is the splittability of the symmetry. In this way, the absence of free parameters and the absence of global symmetry in gravitational theories in AdS are unified and understood in a single framework as consequences of the splittability of symmetry generators.
Figure 1: Interface CFT. Following standard convention, we denote the stress tensors of the CFTs on the left and right sides of the interface by $T_{L}$ and $T_{R}$, respectively. Although many quantities associated with the interface are labeled by $L$ and $R$, as indicated in the left figure, it is more convenient for explicit computations to place the interface along the real axis in the complex plane, as shown in the right figure, thereby reducing the problem to a boundary CFT setup. Following this convention, we will use the coordinate system depicted in the right figure.
@@
In what follows, we show that exactly marginal operators exist, under the assumption of the existence of a conformal interface between nearby CFTs that reduces to the identity when the CFTs coincide. The argument is constructive and model-independent. We focus on two-dimensional CFTs for simplicity, but the proof can be generalized to higher dimensions and will be reported in [Komatsu].
@@
To motivate our proposal, consider a theory with an exactly marginal operator $\phi$, which we can use to deform the CFT in the left half-plane as (See FIG. 1 for our choice of convention)
@@
where $S_{0}$ is the original action of the CFT. At the leading order in $\delta\lambda$, this defines a conformal interface satisfying the gluing condition at ${\rm Im}\,z=0$ [Oshikawa1996, Oshikawa1997, Bachas2002],
$T_{L}-\overline{T}_{L}=T_{R}-\overline{T}_{R},$ (2)
where $T_{L,R}$ ($\overline{T}_{L,R}$) are the holomorphic (anti-holomorphic) stress tensors on the left and right half planes. Furthermore, it follows from the Ward identity that
$T_{L}(x+i0)-T_{R}(x-i0)=\pi\delta\lambda\phi(x)+O(\delta\lambda^{2})\,$, (3)
where we have written $z=x+iy$. Now we consider the displacement operator associated with this conformal interface,
$\mathrm{D}\equiv T_{L}+\overline{T}_{L}-T_{R}-\overline{T}_{R}=2(T_{L}-T_{R}).$ (4)
@@
where $\mathrm{N}_{\mathrm{D}}$ can be computed using (2) and (3) as [Brunner2016]
@@
In the rest of this letter, we show that one can reverse this logic; assuming the existence of a conformal interface between nearby CFTs that reduces to the identity when the CFTs coincide, we provide a simple proof for the existence of an exactly marginal operator. It is desirable to prove the result without assuming the existence of such an interface. We are actively investigating this and hope to present details in an extended version forthcoming [Komatsu].
@@ II Displacement Operator as Exactly Marginal Operator
Our proof assumes
@@
    For any pair of CFTs that are close to each other on the conformal manifold, there is a conformal interface that becomes trivial when the two CFTs coincide.
@@
The consequence of these two assumptions is that defect operators become local operators of the CFT in the limit, unless they decouple (see FIG. 2). [^3] If the interface is not trivial but topological and not invertible, then a topological line might be attached to the defect operators in the limit. For the reasons explained later, we also make the additional technical assumption that
@@
On the other hand, we do not need to assume that the central charges on both sides ($c_{L,R}$) are the same. Furthermore, chiral and antichiral central charges are allowed to differ, while the condition $c_{L}-\bar{c}_{L}=c_{R}-\bar{c}_{R}$ is necessary for the existence of a conformal interface [Billo:2016cpy]. In other words, the central charge can vary on the conformal manifold, but $c-\bar{c}$ cannot.  [^4] From modular $T$-invariance, $h-\bar{h}-\frac{c-\bar{c}}{24}\in\mathbb{Z},$ (9) while the single-valuedness of correlation functions implies the quantization of spin, $h-\bar{h}\in\mathbb{Z}/2$. Combining these relations, we conclude that $c-\bar{c}$ must be quantized. This, in turn, means that $c-\bar{c}$ is invariant under continuous deformations.
@@
Let us show that the displacement operator $\hat{\mathrm{D}}$ becomes an interface primary of weight $2$ in the limit where the two CFTs coincide. When evaluated on the interface, the stress tensors reorganize in two interface primaries ($W$, $\bar{W}$) and a quasiprimary ($T_{tot}=T_{L}+\overline{T}_{R}$), see [Quella2007]. Notice that the gluing condition implies $T_{tot}=\overline{T}_{tot}$. In particular, the displacement operator $\rm D$} can be shown to be an interface quasiprimary. Indeed, one can check that
$\left(T_{L}(z)+\overline{T}_{R}(z)\right)\hat{\mathrm{D}}(w)=\frac{(c_{L}-c_{R})/\sqrt{\mathrm{N}_{\mathrm{D}}}}{(z-w)^{4}}$ (10)
@@
where $z$ and $w$ lie on the real axis, and $O(1)$ denotes terms that are non-singular at $z=w$. Equation (10) is obtained by making use of the interface boundary condition (2), as well as of the fact that $T_{L}(z)\overline{T}_{R}(w)$ and $\overline{T}_{L}(z)T_{R}(w)$ have no singularities at $z=w$. This absence of singularity follows from the cluster decomposition principle [Billo:2016cpy].
In the limit ${\rm N}_{\rm D}\to 0$, the singular part of (10) vanishes, thanks to the Averaged Null Energy conditions, which enforce the bound $\mathrm{N}_{\mathrm{D}}\geq 2\left|c_{L}-c_{R}\right|$ [Meineri2020]. [^5] Reflection positivity is only sufficient to show that (2) is bounded in the limit [Billo:2016cpy], while the smoothness condition implies the even stronger bound; $\left|c_{L}-c_{R}\right|^{2/3}/{\rm N}_{\mathrm{D}}$ stays finite in the limit ${\rm N}_{\rm D}\to 0$, see the Supplemental Material. We obtain the OPE
@@
Given that $\hat{\mathrm{D}}$ is an interface operator of conformal dimension 2, this constrains the possible bulk local operators that can emerge in the ${\rm N}_{\rm D}\to 0$ limit. In the following, we examine each candidate and show that the only viable possibility is a primary operator with conformal dimensions $(1,1)$. We broadly distinguish two cases: a highest weight state under the conformal symmetry $\mathrm{Vir}\otimes\overline{\mathrm{Vir}}$ (i.e., a primary state), and those that can be obtained by acting with the stress tensor modes $L_{-n}$ on primary states (i.e., descendant states).
@@
Recalling that $\phi$ is constructed from $T_{L}(w)-T_{R}(w)$, this expression implies that the OPE between $T_{L}(z)-\overline{T}_{L}(z)$ and $T_{L}(w)-T_{R}(w)$ contains symmetric singular terms, unless $h-\bar{h}=0$. On the other hand, the interface boundary conditions on the energy-momentum tensor imply that the singular parts of the OPE are anti-symmetric in $z$ and $w$,
$\left(T_{L}(z)-\overline{T}_{L}(z)\right)\left(T_{L}(w)-T_{R}(w)\right)$ (15)
$=T_{L}(z)\overline{T}_{L}(w)-\overline{T}_{L}(z)T_{L}(w)+O(1),$
where we used the fact that $T_{L}(z)\overline{T}_{R}(w)$ and $\overline{T}_{L}(z)T_{R}(w)$ do not have singularities at $z=w$. Consequently, we must have $h=\bar{h}$. Since $h+\bar{h}=2$, the bulk primary field $\phi$ that can appear in $\hat{\mathrm{D}}$ must be of conformal weights $(1,1)$, i.e., marginal.
@@
Let us now verify that the three-point function of $\phi$ vanishes, as expected from exact marginality. Given a pair of quasiprimary operators $O_{L}$ and $O_{R}$, with the same spin $h_{L}-\bar{h}_{L}=h_{R}-\bar{h}_{R}$, on either side of any conformal interface, their correlators with and without the displacement are constrained as follows [Gliozzi:2015qsa]:
$\braket{O_{L}(z,\bar{z})O_{R}(\bar{z},z)}=\frac{a_{LR}}{\left(2\mathrm{Im}z\right)^{\Delta_{L}+\Delta_{R}}}$ (21)
$\braket{O_{L}(z,\bar{z})O_{R}(\bar{z},z)\hat{\rm D}(w)}=\frac{\left(\Delta_{L}-\Delta_{R}\right)a_{LR}/\sqrt{\mathrm{N}_{\mathrm{D}}}}{\left(2\mathrm{Im}z\right)^{\Delta_{L}+\Delta_{R}-2}|z-w|^{4}}~,$ (22)
where $\Delta=h+\bar{h}$, $w\in\mathbb{R}$, $a_{LR}$ is a constant which, at $\mathrm{N}_{\mathrm{D}}=0$, is fixed by normalization, i.e., $a_{LR}=1$, and it should be noted that $O_{L}$ and $O_{R}$ are placed at symmetric positions with respect to the interface. Having established the existence of an operator of weight 2 at a generic point of the conformal manifold, we select $O_{L}=O_{R}=\phi$, so that the r.h.s. of (22) vanishes identically for any value of $\mathrm{N}_{\mathrm{D}}$, while (21) does not, and in particular $\lim_{\mathrm{N}_{\mathrm{D}}\to 0}a_{LR}=1$. Since the r.h.s. of (22) reduces to the three-point function of $\phi$ at $\mathrm{N}_{\mathrm{D}}=0$, we conclude that
@@ III Interface Conformal Manifold
Why did we need to assume the absence of neutral $(\frac{1}{2},\frac{1}{2})$ primaries? Let us first clarify what $\rho$ actually represents in (20). If a CFT contains a weight one scalar operator, one can produce a line defect by integrating it with an infinitesimal coefficient. If, furthermore, $c_{\rho\rho\rho}=0$, together with the required higher order conditions, the line defect is part of a conformal manifold. As a concrete example, the Ising CFT admits a defect conformal manifold generated by the following interface exactly marginal deformation [Oshikawa1997],
@@
This is precisely of the same form as the second term in (20). Vice versa, given a family of interfaces connected to the identity, whose displacement obeys (20) with $\phi=0$, one can prove that $c_{\rho\rho\rho}=0$—see the Supplemental Material. Furthermore, it is easy to check that, if $\lim_{\mathrm{N}_{\mathrm{D}}\to 0}\hat{\rm D}=(\partial-\bar{\partial})\rho$, the l.h.s. of (22) vanishes in the limit, therefore no scaling dimension changes at leading order in $\mathrm{N}_{\mathrm{D}}.$ More generally, symmetry-neutral exactly marginal deformation on the interface can arise from the breaking of a phantom symmetry [Antinucci2025]: a symmetry that emerges when one interprets the interface between $\mathrm{CFT}_{L}$ and $\mathrm{CFT}_{R}$ as a boundary in $\mathrm{CFT}_{L}\otimes\overline{\mathrm{CFT}_{R}}$ using the folding trick. In such cases, a phantom current with conformal dimension $(1,0)$ in the folded theory induces an interface exactly marginal operator with dimension $(1/2,1/2)$ in the unfolded theory. This is precisely the case in the example of the Ising CFT, where the phantom current is given by $j=\psi_{L}\psi_{R}$. Here, $\psi_{L/R}$ represents the fermion field on the left/right side of the interface, and the energy density operator in the Ising CFT can be expressed in terms of these fermions as $\varepsilon_{L/R}=i\psi_{L/R}\bar{\psi}_{L/R}$.

### v2

@@ I Introduction and Summary
If exactly marginal operators exist in a conformal field theory (CFT), they generate a conformal manifold, which parameterizes a continuous family of CFTs [29]. One might naturally ask if the converse also holds: Does a conformal manifold require the existence of exactly marginal operators? In this letter, we address this question by deriving the existence of exactly marginal operators under certain assumptions.
@@
The existence of exactly marginal operators has important implications for the Swampland program, particularly for the Distance Conjecture [18]. The earliest distance conjecture may have been stated by Albert Einstein in his autobiographical notes, in which he claimed that there are no free parameters in the basic equations of physics. This was reformulated as Conjecture 0 in [18] as “the moduli space of a consistent quantum gravity is parameterized by inequivalent expectation values of massless scalar fields.” The existence of exactly marginal operators for a conformal manifold allows us to prove Conjecture 0 for gravitational theories in anti-de Sitter (AdS) space as follows [20]: If there is a continuous family of gravitational theories in AdS, then there must be a corresponding family of CFTs. If we can show that each CFT parameter is associated with an exactly marginal operator, the AdS/CFT dictionary would tell us that the exactly marginal operator corresponds to a massless scalar field with a flat potential, and the asymptotic behavior of this field determines the parameters. Therefore, every parameter in quantum gravity in AdS is related to a value of a dynamical field, as claimed in Conjecture 0.111There has been an interesting development in interpreting the conjecture in AdS [6, 25, 27, 26].
@@
We note an interesting analogy with the absence of global symmetry in quantum gravity in AdS, which was proven in [12, 13] using the splittability of symmetry generators [10] in CFT. For continuous symmetry, the splittability is the consequence of the well-known fact that the symmetry generators are expressed as exponentiated integrals of the Noether currents over a Cauchy surface. If we regard a continuous deformation of CFT as (1)(-1)-form global symmetry, the existence of the corresponding exactly marginal operator is the splittability of the symmetry. In this way, the absence of free parameters and the absence of global symmetry in gravitational theories in AdS are unified and understood in a single framework as consequences of the splittability of symmetry generators.
Figure 1: Interface CFT. We denote the stress tensors of the CFTs on the upper and bottom sides of the interface by $T_{1}$ and $T_{2}$, respectively.
@@
In what follows, we show that exactly marginal operators exist, under the assumption of the existence of a conformal interface between nearby CFTs that reduces to the identity when the CFTs coincide. The argument is constructive and model-independent. We focus on two-dimensional CFTs for simplicity, but the proof can be generalized to higher dimensions and will be reported in [15].
@@
To motivate our proposal, consider a theory with an exactly marginal operator $\phi$, which we can use to deform the CFT in the upper half-plane as (See FIG. 1 for our choice of convention)
@@
where $S_{0}$ is the original action of the CFT. At the leading order in $\delta\lambda$, this defines a conformal interface satisfying the gluing condition at ${\rm Im}\,z=0$ [21, 22, 5],
$T_{1}-\overline{T}_{1}=T_{2}-\overline{T}_{2},$ (2)
where $T_{1,2}$ ($\overline{T}_{1,2}$) are the holomorphic (anti-holomorphic) stress tensors on the upper and bottom half planes. Furthermore, it follows from the Ward identity that
$T_{1}(x+i0)-T_{2}(x-i0)=\pi\delta\lambda\phi(x)+O(\delta\lambda^{2})\,$, (3)
where we have written $z=x+iy$. Now we consider the displacement operator associated with this conformal interface [8],
$\mathrm{D}\equiv T_{1}+\overline{T}_{1}-T_{2}-\overline{T}_{2}=2(T_{1}-T_{2}).$ (4)
@@
where $\mathrm{N}_{\mathrm{D}}$ can be computed using (2) and (3) as [9]
@@
In the rest of this letter, we show that one can reverse this logic; assuming the existence of a conformal interface between nearby CFTs that reduces to the identity when the CFTs coincide, we provide a simple proof for the existence of an exactly marginal operator. It is desirable to prove the result without assuming the existence of such an interface. We are actively investigating this and hope to present details in an extended version forthcoming [15].
@@ II Displacement Operator as Exactly Marginal Operator
We define a conformal manifold as a collection of CFT data which depend continuously on a set of parameters. For every choice of parameters, the stress tensor exists and generates two copies of the Virasoro algebra. Our proof further assumes that
@@
    For any pair of CFTs that are close to each other on the conformal manifold, there is a conformal interface that becomes trivial (i.e. the identity interface) when the two CFTs coincide.
@@
The consequence of these two assumptions is that defect operators become local operators of the CFT in the limit, unless they decouple (see FIG. 2). [^3] If the interface in the limit is not trivial but topological and non-invertible, defect operators may become operators with a topological line attached. This is true, in particular, for the (normalized) displacement operator $\hat{\mathrm{D}}$, which exists on every conformal interface. For the reasons explained later, we also make the additional technical assumption that
@@
Note that this third assumption excludes Liouville theory from our discussion [^4] For arbitrary values of the coupling constant, Liouville theory has a symmetry-neutral $(1/2,1/2)$ operator, hence violating our assumption 3. This operator may not correspond to a normalizable state in Liouville theory, but our current argument is insensitive to such a normalizability issue..
On the other hand, we do not need to assume that the central charges on both sides ($c_{L,R}$) are the same. Furthermore, chiral and antichiral central charges are allowed to differ, while the condition $c_{1}-\bar{c}_{1}=c_{2}-\bar{c}_{2}$ is necessary for the existence of a conformal interface [8]. In other words, the central charge can vary on the conformal manifold, but $c-\bar{c}$ cannot. [^5] From modular $T$-invariance, $h-\bar{h}-\frac{c-\bar{c}}{24}\in\mathbb{Z},$ (9) while the single-valuedness of correlation functions implies the quantization of spin, $h-\bar{h}\in\mathbb{Z}/2$. Combining these relations, we conclude that $c-\bar{c}$ must be quantized. This, in turn, means that $c-\bar{c}$ is invariant under continuous deformations.
@@
Let us show that the displacement operator $\hat{\mathrm{D}}$ becomes an interface primary of weight $2$ in the limit where the two CFTs coincide. When evaluated on the interface, the stress tensors reorganize in two interface primaries ($W$, $\bar{W}$) and a quasiprimary ($T_{tot}=T_{1}+\overline{T}_{2}$), see [24]. Notice that the gluing condition implies $T_{tot}=\overline{T}_{tot}$. In particular, the displacement operator $\rm D$} can be shown to be an interface quasiprimary. Indeed, one can check that
$\left(T_{1}(z)+\overline{T}_{2}(z)\right)\hat{\mathrm{D}}(w)=\frac{(c_{1}-c_{2})/\sqrt{\mathrm{N}_{\mathrm{D}}}}{(z-w)^{4}}$ (10)
@@
where $z$ and $w$ lie on the real axis, and $O(1)$ denotes terms that are non-singular at $z=w$. Equation (10) is obtained by making use of the interface boundary condition (2), as well as of the fact that $T_{1}(z)\overline{T}_{2}(w)$ and $\overline{T}_{1}(z)T_{2}(w)$ have no singularities at $z=w$. This absence of singularity follows from the cluster decomposition principle [8].
In the limit ${\rm N}_{\rm D}\to 0$, the singular part of (10) vanishes, thanks to the Averaged Null Energy conditions, which enforce the bound $\mathrm{N}_{\mathrm{D}}\geq 2\left|c_{1}-c_{2}\right|$ [16]. [^6] Reflection positivity is only sufficient to show that (2) is bounded in the limit [8], while the smoothness condition implies the even stronger bound; $\left|c_{1}-c_{2}\right|^{2/3}/{\rm N}_{\mathrm{D}}$ stays finite in the limit ${\rm N}_{\rm D}\to 0$, see the Supplemental Material. We obtain the OPE
@@
Given that $\hat{\mathrm{D}}$ is an interface operator of conformal dimension 2, this constrains the possible bulk local operators that can emerge in the ${\rm N}_{\rm D}\to 0$ limit. In the following, we examine each candidate and show that the only viable possibility is a primary operator with conformal dimensions $(1,1)$. We broadly distinguish two cases: a highest weight state under the conformal symmetry $\mathrm{Vir}\otimes\overline{\mathrm{Vir}}$ (i.e. a primary state), and those that can be obtained by acting with the stress tensor modes $L_{-n}$ on primary states (i.e. descendant states).
@@
Recalling that $\phi$ is constructed from $T_{1}(w)-T_{2}(w)$, this expression implies that the OPE between $T_{1}(z)-\overline{T}_{1}(z)$ and $T_{1}(w)-T_{2}(w)$ contains symmetric singular terms, unless $h-\bar{h}=0$. On the other hand, the interface boundary conditions on the energy-momentum tensor imply that the singular parts of the OPE are anti-symmetric in $z$ and $w$,
$\left(T_{1}(z)-\overline{T}_{1}(z)\right)\left(T_{1}(w)-T_{2}(w)\right)$ (15)
$=T_{1}(z)\overline{T}_{1}(w)-\overline{T}_{1}(z)T_{1}(w)+O(1),$
where we used the fact that $T_{1}(z)\overline{T}_{2}(w)$ and $\overline{T}_{1}(z)T_{2}(w)$ do not have singularities at $z=w$. Consequently, we must have $h=\bar{h}$. Since $h+\bar{h}=2$, the bulk primary field $\phi$ that can appear in $\hat{\mathrm{D}}$ must be of conformal weights $(1,1)$, i.e., marginal.
@@
Let us now verify that the three-point function of $\phi$ vanishes, as expected from exact marginality. Given a pair of quasiprimary operators $O_{1}$ and $O_{2}$, with the same spin $h_{1}-\bar{h}_{1}=h_{2}-\bar{h}_{2}$, on either side of any conformal interface, their correlators with and without the displacement are constrained as follows [11]:
$\braket{O_{1}(z,\bar{z})O_{2}(\bar{z},z)}=\frac{a_{12}}{\left(2\mathrm{Im}z\right)^{\Delta_{1}+\Delta_{2}}}$ (21)
$\braket{O_{1}(z,\bar{z})O_{2}(\bar{z},z)\hat{\rm D}(w)}=\frac{\left(\Delta_{1}-\Delta_{2}\right)a_{12}/\sqrt{\mathrm{N}_{\mathrm{D}}}}{\left(2\mathrm{Im}z\right)^{\Delta_{1}+\Delta_{2}-2}|z-w|^{4}}~,$ (22)
where $\Delta=h+\bar{h}$, $w\in\mathbb{R}$, $a_{12}$ is a constant which, at $\mathrm{N}_{\mathrm{D}}=0$, is fixed by normalization, i.e., $a_{12}=1$, and it should be noted that $O_{1}$ and $O_{2}$ are placed at symmetric positions with respect to the interface. Having established the existence of an operator of weight 2 at a generic point of the conformal manifold, we select $O_{1}=O_{2}=\phi$, so that the r.h.s. of (22) vanishes identically for any value of $\mathrm{N}_{\mathrm{D}}$, while (21) does not, and in particular $\lim_{\mathrm{N}_{\mathrm{D}}\to 0}a_{12}=1$. Since the r.h.s. of (22) reduces to the three-point function of $\phi$ at $\mathrm{N}_{\mathrm{D}}=0$, we conclude that
@@ III Interface Conformal Manifold
Why did we need to assume the absence of neutral $(\frac{1}{2},\frac{1}{2})$ primaries? Let us first clarify what $\rho$ actually represents in (20). If a CFT contains a weight one scalar operator, one can produce a line defect by integrating it with an infinitesimal coefficient. If, furthermore, $c_{\rho\rho\rho}=0$, together with the required higher order conditions, the line defect is part of a conformal manifold. As a concrete example, the Ising CFT admits a defect conformal manifold generated by the following interface exactly marginal deformation [22],
@@
This is precisely of the same form as the second term in (20). Vice versa, given a family of interfaces connected to the identity, whose displacement obeys (20) with $\phi=0$, one can prove that $c_{\rho\rho\rho}=0$—see the Supplemental Material. Furthermore, it is easy to check that, if $\lim_{\mathrm{N}_{\mathrm{D}}\to 0}\hat{\rm D}=(\partial-\bar{\partial})\rho$, the l.h.s. of (22) vanishes in the limit, therefore no scaling dimension changes at leading order in $\mathrm{N}_{\mathrm{D}}.$ More generally, symmetry-neutral exactly marginal deformation on the interface can arise from the breaking of a phantom symmetry [3]: a symmetry that emerges when one interprets the interface between $\mathrm{CFT}_{1}$ and $\mathrm{CFT}_{2}$ as a boundary in $\mathrm{CFT}_{1}\otimes\overline{\mathrm{CFT}_{2}}$ using the folding trick. In such cases, a phantom current with conformal dimension $(1,0)$ in the folded theory induces an interface exactly marginal operator with dimension $(1/2,1/2)$ in the unfolded theory. This is precisely the case in the example of the Ising CFT, where the phantom current is given by $j=\psi_{1}\psi_{2}$. Here, $\psi_{1/2}$ represents the fermion field on the upper/bottom side of the interface, and the energy density operator in the Ising CFT can be expressed in terms of these fermions as $\varepsilon_{1/2}=i\psi_{1/2}\bar{\psi}_{1/2}$.